OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 2x+5y=16 -5x-2y=2

OpenStudy (issy14):

2x+5=16 -5x-2y=2 2x=16-5 -5(5.5)-2y=2 2x=11 -27.5-2y=2 x=11/2 -2y=2+27.5 x= 5.5 -2y=29.5 y=29.5/-2 y=-14.75 ordered pair: (5.5,-14.75)